# A Survey of Multimodal Large Language Model from A Data-centric Perspective

### v1

@@ 3.3.3. Audio Datasets
Audio datasets are curated collections of sound recordings along with associated annotations or labels. These datasets are specifically designed to support the training and evaluation of models for various audio-related tasks, such as speech recognition, sound event detection, music classification, and speaker identification. Audio datasets vary significantly in size, quality, and scope, with some focusing on specific languages or acoustic environments, while others strive for diversity to ensure generalizability across different audio processing scenarios. Notable examples include AISHELL-2 (Du et al., 2018), a Mandarin Chinese speech corpus with over 1000 hours of data from multiple regions; WavCaps (Mei et al., 2023), an extensive English audio captioning dataset featuring approximately 400,000 clips; and VSDial-CN (Chen et al., 2023b), a multi-modal dataset derived from VisDial, encompassing visual data alongside related dialogues and captions, tailored for Automatic Speech Recognition (ASR) systems. These datasets are instrumental in advancing the capabilities of language models, enhancing their ability to understand and generate human-like responses in various audio-visual contexts.
@@
The development of MLLMs benefits from the utilization of diverse 3D datasets, which provide comprehensive environmental and structural data critical for tasks like scene understanding and semantic segmentation. Notably, the ScanNet (Dai et al., 2017) dataset offers an extensive collection of RGB-D video data across 1,513 indoor scenes, annotated with 3D camera poses, surface reconstructions, and semantic segmentations, totaling over 2.5 million views. Similarly, the S3DIS (Armeni et al., 2017) dataset includes point clouds from six large indoor areas, encompassing 271 rooms, with detailed semantic annotations for each point. The Structured3D (Zheng et al., 2020a) dataset provides a vast repository of 3,500 home designs, encompassing 21,835 rooms detailed with elements such as object geometry, materials, and textures, tailored for analyses in 3D reconstruction and interior design. Together, these datasets present a rich framework for training and evaluating the capabilities of LLMs in interpreting and generating three-dimensional spatial data.
@@ 4. Data-centric pre-training
The second phase builds upon the knowledge acquired in the first phase by further training the LLM and modality encoders using a mixture of multimodal data. During this phase, the input projector (such as a linear projector, cross-attention mechanism, Q-Former (Li et al., 2022b), or P-Former (Jian et al., 2024)) is trained to effectively map the features extracted by the modality encoders into the LLM’s embedding space. This mapping allows for a unified representation of textual and non-textual information within the model’s architecture.
Some researchers consider employing a selective training approach during the second phase, which involves keeping certain components of the model frozen while training specific parts. This approach helps to preserve the knowledge acquired during the first phase, allows for targeted adaptation to specific multimodal tasks, and reduces computational costs.
@@ 4.1. Domain Mixture
Previous methods typically select domain weights heuristically or optimize them using downstream tasks (Du et al., 2022). However, these approaches can be suboptimal or costly, requiring the training of LMs for different sets of domain weights and potentially leading to overfitting to specific downstream tasks. DoReMi (Xie et al., 2024a) offers a solution by optimizing domain weights using proxy models that are 30 times smaller than the target LLM, without requiring knowledge of downstream tasks.
In the realm of video understanding, there is a consensus on the necessity for diverse datasets (Xu et al., 2023c; Chen et al., 2024b; Song et al., 2023; Chen et al., 2023l; Wang et al., 2023a). Xu et al. (2023c) highlight the use of hierarchical multi-label classification models (Giunchiglia and Lukasiewicz, 2020) to ensure dataset balance and comprehensiveness across multiple dimensions, thereby optimizing the model’s ability to understand and process various video domains. Such diversity is instrumental in training more versatile and universally applicable video-language models.
@@ 4.2. Modality Mixture
When pre-training MLLMs, determining the optimal proportions of multimodal data is crucial for enhancing the model’s performance across different tasks.
@@
MM1 (McKinzie et al., 2024) explores the balance between image-caption pairs, interleaved image-caption documents, and text-only data for vision-language pre-training. They found that a ratio of 5:5:1 for caption/interleaved/text data yields the best overall performance on text-only tasks as well as zero-shot and few-shot image-text tasks. Additionally, incorporating synthetic data (Lai et al., 2023b) improves the model’s few-shot learning capabilities for image-text tasks.
For the development of MLLMs tailored for video understanding, the strategic integration of both video-text and image-text pairs during the pre-training phase is pivotal. This approach, highlighted in recent research (Lin et al., 2023b; Zhang et al., 2023; Luo et al., 2023; Jin et al., 2023; Li et al., 2023g; Han et al., 2023; Li et al., 2023c, h; Wang et al., 2023a; Chen et al., 2023c), enhances the model’s capabilities by leveraging the complementary nature of these data types.
Integrating labeled image datasets like COCO with video data broadens the variety of training datasets and addresses the scarcity of high-quality video-text resources. Studies (Lin et al., 2023b; Jin et al., 2023) indicate that training on both images and videos enhances models’ ability to understand static and dynamic visual information without needing task-specific adaptations. Adding a temporal modeling module to the vision encoder bridges the temporal dynamics of videos with the static nature of images, fostering cohesive visual understanding across different types of visual media.
However, treating images as single-frame videos (Zhang et al., 2023; Luo et al., 2023) or pseudo-videos may weaken the model’s ability to grasp the temporal aspects of video sequences. An imbalanced mix could also bias the model towards one modality, complicating its learning process. Therefore, determining the optimal mix ratio of image-text pairs and video-text pairs is crucial for improving the model’s comprehensive understanding capabilities, ensuring a balanced approach to learning from both static images and dynamic videos.
@@
Additionally, video LLMs are increasingly being trained through multiple branches (Chen et al., 2024b; Zhang et al., 2023; Lyu et al., 2023; Shu et al., 2023b; Han et al., 2023; Sun et al., 2023c; Chen et al., 2023c), such as vision-language, audio-language, and subtitle-language. These methods leverage the correlations among different modalities to broaden the models’ understanding and learning capabilities. Recent studies (Shu et al., 2023b; Chen et al., 2024b; Han et al., 2023) have shown that combining audio, subtitles, and visual data during training enhances performance across various video understanding benchmarks. This modular training approach also offers the flexibility to pre-train with partial data (Han et al., 2023; Sun et al., 2023c), such as visual-only or audio-only, particularly when comprehensive multimodal data is not available.
@@ 4.3. Quality Selection
Due to the diverse distribution of data, training a large model with all available data is not always optimal. Therefore, data selection becomes essential, offering benefits such as reduced training time and energy consumption (Gadre et al., 2024). Unlike pure text datasets, data selection for multimodal datasets must consider the alignment between different modalities, in addition to metrics like perplexity used for single-modality evaluation (Wang et al., 2024d).
Data selection methods can be categorized into two types: active learning-based and pre-training selection. Active learning-based methods, such as CiT (Xu et al., 2023a), use data proxies to dynamically select training data during the training process, achieving significant acceleration effects. Pre-training selection methods, like Datacomp (Gadre et al., 2024), evaluate and select all data before training begins.
Regarding data selection criteria, methods can be distribution-agnostic, focusing solely on individual data point quality, or distribution-aware, considering the overall data distribution. Distribution-agnostic methods include training a model with the top 30% of data ranked by CLIP score, which can significantly improve results (Gadre et al., 2024). More comprehensive metrics have also been developed, showing better performance than using CLIP score alone (Wang et al., 2024d). For instance, Mahmoud et al. (2024) use the difference between original and synthetic captions generated by a small model to evaluate image-text pair alignment.
Given the limited scope of purely distribution-agnostic work, combining distribution-agnostic and distribution-aware methods is often more effective. For example, using CLIP-score-based filters alongside image-based filters can outperform either method alone on large datasets (Wang et al., 2024a; Gadre et al., 2024).
@@
Adaptation is crucial for aligning pre-trained multimodal large language models (MLLMs) with specific tasks and user preferences.
Self-supervised pre-training provides LLMs with a broad understanding of textual and multimodal information, while supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) are essential for adapting these models to excel in targeted applications and adhere to user preferences and societal values.
@@
Data-centric SFT involves training models on carefully curated datasets across multiple modalities. In contrast, data-centric RLHF focuses on collecting human judgments to guide the model toward generating desirable and ethically aligned responses. Emphasizing data-centric approaches highlights the importance of high-quality, domain-specific, and multimodal datasets in the successful adaptation of MLLMs.
Table 1. A detailed list of datasets and processing methods for different models with image modality during fine-tuning stage. * indicates the dataset is newly generated using certain method within a respective model, while the other datasets (without *) serve as the original data sources for that model. - denotes directly using original data sources without additional processing. #Examples denotes the statistics for each dataset.
Models Time Finetuning Datasets Processing Method #Examples
Flamingo(Alayrac et al., 2022) 2022.4 VQAv2 - #images: 265K    #questions: 1.4M
VizWiz #image-question pairs: 33.8K
TextVQA #images: 28.4K    #questions: 45.3K
BLIP-2(Li et al., 2023f) 2023.01 COCO - #images: 330K    #captions: 5 per image
NoCaps #images: 15.1K    #captions: 11 per image
VQAv2 #images: 265K    #questions: 1.4M
Flickr30K #images: 31.7K    #captions: 5 per image
LLaVA(Liu et al., 2024b) 2023.04 COCO GPT-4, Prompt, Manual examples #images: 330K    #captions: 5 per image
LLaVA-Instruct-150K* #image-text pairs: 158K
MiniGPT-4(Zhu et al., 2023) 2023.04 CC12M Manual verifying and refining, ChatGPT, Random selecting #image-text pairs: 12M
cc_sbu_align* #image-text pairs: 3.5K
mPLUG-Owl(Ye et al., 2023b) 2023.04 LLaVA-Instruct-150K - #image-text pairs: 158K
InstructBLIP(Dai et al., 2024) 2023.05 ScienceQA - #image-question pairs: 10.3K
OCR-VQA #images: 207K    #image-question pairs: 1M
OKVQA #images: 14K    #questions: 14K
A-OKVQA #images: 23.6K    #questions: 24.9K
PaLI-X(Chen et al., 2023a) 2023.05 COCO - #images: 330K    #captions: 5 per image
NoCaps #images: 15.1K    #captions: 11 per image
TextCaps #images: 28.4K    #captions: 5 per image
VizWizCap #images: 39.1K    #captions: 5 per image
Screen2Words #images: 22.4K    #captions: 5 per image
WidgetCap #images: 21.7K    #captions: 162K    #widgets: 61.2K
Shikra(Chen et al., 2023j) 2023.06 LLaVA-Instruct-150K GPT-4, Prompt, Sampling ratio: 0.5 #image-text pairs: 158K
Flickr30K #images: 31.7K    #captions: 5 per image
Shikra-RD* #qustion-answer pairs: 5.9K
DLP(Jian et al., 2024) 2023.07 COCO - #images: 330K    #captions: 5 per image
ChatSpot(Zhao et al., 2023c) 2023.07 Visual Genome Unification, GPT-4, Prompt #images: 108K    #region descriptions: 5.4M    #image-question pairs: 1.7M
MiniGPT-5(Zheng et al., 2023) 2023.10 VIST Flexible framework for various task, Placeholders for images -¿ Text prompts #images: 210K    #stories: 50K
MMDialog #images: 1.53M    #texts: 1.08M    #turns: 4.92M
LLaVA-1.5(Liu et al., 2023c) 2023.10 RefCOCO Merging and concatenating, Splitting, Sampling #images: 20K    #captions: 142K
GQA #images: 113K    #questions: 22M
OCR-VQA #images: 207K    #image-question pairs: 1M
TextCaps #images: 28.4K    #captions: 5 per image
Visual Genome #images: 108K    #region descriptions: 5.4M    #image-question pairs: 1.7M
llava_v1_5_mix665k* #image-text pairs: 665K
MiniGPT-v2(Chen et al., 2023m) 2023.10 LLaVA-Instruct-150K Object parsing and grounding, Selecting captions #image-text pairs: 158K
Flickr30K #images: 31.7K    #captions: 5 per image
CogVLM(Wang et al., 2024c) 2023.11 LLaVA-Instruct-150K Translation into Chinese, Correcting and retranslation #image-text pairs: 158K
cc_sbu_align #image-text pairs: 3.5K
CogVLM-SFT-311K* #images: 155K    #image-text pairs: 311K
DRESS(Chen et al., 2023i) 2023.11 LLaVA-Instruct-150K Partition the multi-turn into separate turns, LLM filtering, LLM-Human-in-the-Loop process #image-text pairs: 158K
COCO #images: 330K    #captions: 5 per image
VLSafe* #image-text pairs: 5.8K
VILA(Lin et al., 2023a) 2023.12 llava_v1_5_mix665k - #image-text pairs: 665K
ShareGPT4V(Chen et al., 2023e) 2023.11 COCO GPT-4, Data-specific prompt, Data replacement #images: 330K    #captions: 5 per image
LAION2B-en #image-text pairs: 2.32B
CC3M #image-text pairs: 3.3M
SBU Captions #image-text pairs: 1M
SAM #images: 11M    #segmentation masks: 1.1B
TextCaps #images: 28.4K    #captions: 5 per image
WikiArt #images: 81.4K    #artist classes: 129    #genre classes: 11    #style classes: 27
llava_v1_5_mix665k #image-text pairs: 665K
ShareGPT4V-1.2M* #image-text pairs: 1.2M
GLaMM(Rasheed et al., 2023) 2023.11 RefCOCO Re-purposing of data sources, Manually annotation #images: 20K    #captions: 142K
RefCOCO+ #images: 20K    #captions: 141K
RefCOCOg #images: 25.8K    #captions: 95K
Visual Genome #images: 108K    #region descriptions: 5.4M    #image-question pairs: 1.7M
LLaVA-Instruct-150K #image-text pairs: 158K
GranD-f* #image-text pairs: 214K
PVIT(Chen et al., 2023h) 2023.08 GQA ChatGPT, Prompt, Multi-turn data #images: 113K    #questions: 22M
VCR #images: 110K    #questions: 290K
COCO #images: 330K    #captions: 5 per image
Visual Genome #images: 108K    #region descriptions: 5.4M    #image-question pairs: 1.7M
COCO-Text #images:63.6K    #labeled text regions: 173K
PVIT* #image-text pairs: 13.7M (Stage 1)    #image-question pairs: 10.4K (Stage 2)
TextMonkey(Liu et al., 2024d) 2024.03 COCO-Text Structured data (documents, tables, charts), 5% of pre-train data, #images:63.6K    #labeled text regions: 173K
TextOCR #images:28.1K    #texts: 903K
HierText #images:11.6K    #texts: 1.2M
TextVQA #images: 28.4K    #questions: 45.3K
MLT #image-text pairs: 10K
ChartQA #image-text pairs: 20.8K    #tables: 20.8K
DocVQA #images: 12K    #questions: 50K
InfoVQA #images: 5K    #questions: 30K
Monkey_Data* #image-text pairs: 409.1K
Table 2. A detailed list of datasets and processing methods for different models with video modality during fine-tuning stage. * indicates the dataset is newly generated using certain method within a respective model, while the other datasets (without *) serve as the original data sources for that model. - denotes directly using original data sources without additional processing. #Examples denotes the statistics for each dataset.
Models Time Finetuning Datasets Processing Method #Examples
VideoChat(Li et al., 2023c) 2023.05 WebVid GPT-4, Prompt, Randomly choosing #video-text pairs: 10M
VideoChat* #video-description pairs: 7K    #video-conversation pairs: 4K
Video-LLaMA(Lin et al., 2023b) 2023.01 VideoChat - #video-description pairs: 7K    #video-conversation pairs: 4K
Video-ChatGPT(Maaz et al., 2023) 2023.01 ActivityNet Human-assisted annotations, GPT, Semi-automatic annotation #videos: 20K    #texts: 100K
VideoInstruct100K* #video-instruction pairs: 100K
Table 3. A detailed list of datasets and processing methods for different models with audio modality during the fine-tuning stage. * indicates the dataset is newly generated using certain methods within a respective model, while the other datasets (without *) serve as the original data sources for that model. The datasets in other modals used for generating multi-modal datasets (including audio) are also shown in this table. - denotes directly using original data sources without additional processing. #Examples denote the statistics for each dataset.
Models Time Finetuning Datasets Processing Method #Examples
BuboGPT(Zhao et al., 2023a) 2023.07 Clotho GPT-4, Prompt #audios: 5K    #captions: 24K
VGGSS #video-audio pairs: 5K
Clotho-Detail* #audio-caption pairs: 3.9K    #tokens: 207K
VGGSS-Instruction-Tuning* #audio-image-caption pairs: 5.1K
X-LLM(Chen et al., 2023b) 2023.05 AISHELL-2 Manually selecting, ChatGPT #reading-speech audios: 1000 hours
VSDial-CN #automatic speech recognition (ASR) samples: 1.2M
cc_sbu_align (Image-Text) #image-text pairs: 3.5K
ActivityNet (Video-Text) #videos: 20K    #texts: 100K
NExT-GPT(Wu et al., 2023a) 2023.09 WebVid (Video-Text) External resources, GPT-4, Prompt #video-text pairs: 10M
AudioCaps #audio-text pairs: 46K
CC3M (Image-Text) #image-caption pairs: 3.3M
T2M* #images: 4.9K    #videos: 4.9K    #audios: 4.9K    #instances: 14.7K
MosIT* #images: 4K    #videos: 4K    #audios: 4K    #instances: 5K
X-InstructBLIP(Panagopoulou et al., 2023) 2023.11 AudioCaps Automatic generation, Prompt #audio-text pairs: 46K
AudioCapsQA* #audio QA pairs: 25.4K    #unique questoins: 10.8K
DisCRn* #audio-video pairs: 8.8K
Qwen-Audio(Bai et al., 2023a) 2023.11 Self-constructed Manual annotation, GPT-3.5, Data Mixing #audio-text pairs: 20K

### v2

@@ 3.3.3. Audio Datasets
Audio datasets are curated collections of sound recordings along with associated annotations or labels. These datasets are specifically designed to support the training and evaluation of models for various audio-related tasks, such as speech recognition, sound event detection, music classification, and speaker identification. Audio datasets vary significantly in size, quality, and scope, with some focusing on specific languages or acoustic environments, while others strive for diversity to ensure generalizability across different audio processing scenarios. Notable examples include AISHELL-2 (Du et al., 2018), a Mandarin Chinese speech corpus with over 1000 hours of data from multiple regions; WavCaps (Mei et al., 2023), an extensive English audio captioning dataset featuring approximately 400,000 clips; and VSDial-CN (Chen et al., 2023b), a multi-modal dataset derived from VisDial, encompassing visual data alongside related dialogues and captions, tailored for Automatic Speech Recognition (ASR) systems.
@@
The development of MLLMs benefits from the utilization of diverse 3D datasets, which provide comprehensive environmental and structural data critical for tasks like scene understanding and semantic segmentation. Notably, the ScanNet (Dai et al., 2017) dataset offers an extensive collection of RGB-D video data across 1,513 indoor scenes, annotated with 3D camera poses, surface reconstructions, and semantic segmentations, totaling over 2.5 million views. Similarly, the S3DIS (Armeni et al., 2017) dataset includes point clouds from six large indoor areas, encompassing 271 rooms, with detailed semantic annotations for each point. The Structured3D (Zheng et al., 2020) dataset provides a vast repository of 3,500 home designs, encompassing 21,835 rooms detailed with elements such as object geometry, materials, and textures, tailored for analyses in 3D reconstruction and interior design.
@@ 4. Data-centric pre-training
The second phase builds upon the knowledge acquired in the first phase by further training the LLM and modality encoders using a mixture of multimodal data. During this phase, the input projector (such as a linear projector, cross-attention mechanism, Q-Former (Li et al., 2022a), or P-Former (Jian et al., 2024)) is trained to effectively map the features extracted by the modality encoders into the LLM’s embedding space. This mapping allows for a unified representation of textual and non-textual information within the model’s architecture. Some researchers consider employing a selective training approach during the second phase, which involves keeping certain components of the model frozen while training specific parts. This approach helps to preserve the knowledge acquired during the first phase, allows for targeted adaptation to specific multimodal tasks, and reduces computational costs.
@@ 4.1. Domain Mixture
Previous methods typically select domain weights heuristically or optimize them using downstream tasks (Du et al., 2022). However, these approaches can be suboptimal or costly, requiring the training of LMs for different sets of domain weights and potentially leading to overfitting to specific downstream tasks. DoReMi (Xie et al., 2024a) offers a solution by optimizing domain weights using proxy models that are 30 times smaller than the target LLM, without requiring knowledge of downstream tasks. Recent work (Liu et al., 2024e) formulate domain mixture problem as a regression problem. They assume the rank invariance of domain mixture between small models and large models, thus developing domain mixture regression function to predict the optimal domain mixture for model training. There are also some work discuss this problem through domain mixture scaling law. These work extend the original scaling law (Kaplan et al., 2020; Hoffmann et al., 2022) to domain data mixing law (Que et al., 2024; Ye et al., 2024; Ge et al., 2024), discussing the relationship between domain mixture rate and loss through scaling models.
In the realm of video understanding, there is a consensus on the necessity for diverse datasets (Xu et al., 2023c; Chen et al., 2024b; Song et al., 2023; Chen et al., 2023l; Wang et al., 2023a). Xu et al. (2023c) highlight the use of hierarchical multi-label classification models (Giunchiglia and Lukasiewicz, 2020) to ensure dataset balance and comprehensiveness across multiple dimensions, thereby optimizing the model’s ability to understand and process various video domains, which is instrumental in training more versatile applicable video-language models.
@@ 4.2. Modality Mixture
When pre-training MLLMs, determining the optimal proportions of multimodal data is crucial for enhancing the model’s performance across different tasks. MM1 (McKinzie et al., 2024) explores the balance between image-caption pairs, interleaved image-caption documents, and text-only data for vision-language pre-training. They found that a ratio of 5:5:1 for caption/interleaved/text data yields the best overall performance on text-only tasks as well as zero-shot and few-shot image-text tasks. Additionally, incorporating synthetic data (Lai et al., 2023b) improves the model’s few-shot learning capabilities for image-text tasks.
@@
For the development of MLLMs tailored for video understanding, the strategic integration of both video-text and image-text pairs during the pre-training phase is pivotal. This approach, highlighted in recent research (Lin et al., 2023b; Zhang et al., 2023; Jin et al., 2023; Han et al., 2023; Li et al., 2023c; Wang et al., 2023a), enhances the model’s capabilities by leveraging the complementary nature of these data types. Integrating labeled image datasets like COCO with video data broadens the variety of training datasets and addresses the scarcity of high-quality video-text resources. Studies (Lin et al., 2023b; Jin et al., 2023) indicate that training on both images and videos enhances models’ ability to understand static and dynamic visual information without needing task-specific adaptations. Adding a temporal modeling module to the vision encoder bridges the temporal dynamics of videos with the static nature of images, fostering cohesive visual understanding across different types of visual media. However, treating images as single-frame videos (Zhang et al., 2023; Luo et al., 2023) or pseudo-videos may weaken the model’s ability to grasp the temporal aspects of video sequences. An imbalanced mix could also bias the model towards one modality, complicating its learning process. Therefore, determining the optimal mix ratio of image-text pairs and video-text pairs is crucial for improving the model’s comprehensive understanding capabilities, ensuring a balanced approach to learning from both static images and dynamic videos. Additionally, video LLMs are increasingly being trained through multiple branches (Chen et al., 2024b; Zhang et al., 2023; Lyu et al., 2023), such as vision-language, audio-language, and subtitle-language. These methods leverage the correlations among different modalities to broaden the models’ understanding and learning capabilities. Recent studies (Shu et al., 2023; Chen et al., 2024b; Han et al., 2023) have shown that combining audio, subtitles, and visual data during training enhances performance across various video understanding benchmarks. This modular training approach also offers the flexibility to pre-train with partial data (Han et al., 2023; Sun et al., 2023c), such as visual-only or audio-only, particularly when comprehensive multimodal data is not available.
@@ 4.3. Quality Selection
Due to the diverse distribution of data, training a large model with all available data is not always optimal. Therefore, data selection becomes essential, offering benefits such as reduced training time and energy consumption (Gadre et al., 2024). Previous work has studied data selection methods using n-gram similarity with high quality datasets (Xie et al., 2023; Gao et al., 2020; Chowdhery et al., 2023), perplexity (Marion et al., 2023; Wenzek et al., 2020), influence functions (Park et al., 2023) and llm-based classifier (Wettig et al., 2024). Unlike pure text datasets, data selection for multimodal datasets must consider the alignment between different modalities. Data selection methods can be categorized into two types: active learning-based and pre-training selection. Active learning-based methods, such as CiT (Xu et al., 2023a), use data proxies to dynamically select training data during the training process, achieving significant acceleration effects. Pre-training selection methods, like Datacomp (Gadre et al., 2024) and Bunny (He et al., 2024), evaluate and select all data before training begins.
Regarding data selection criteria, methods can be distribution-agnostic, focusing solely on individual data point quality, or distribution-aware, considering the overall data distribution. Distribution-agnostic methods include training a model with the top 30% of data ranked by CLIP score, which can significantly improve results (Gadre et al., 2024). More comprehensive metrics have also been developed, showing better performance than using CLIP score alone (Wang et al., 2024c). For instance, Mahmoud et al. (2024) use the difference between original and synthetic captions generated by a small model to evaluate image-text pair alignment. Given the limited scope of purely distribution-agnostic work, combining distribution-agnostic and distribution-aware methods is often more effective. For example, using CLIP-score-based filters alongside image-based filters can outperform either method alone on large datasets (Wang et al., 2024a; Gadre et al., 2024).
@@
Adaptation is crucial for aligning pre-trained multimodal large language models (MLLMs) with specific tasks and user preferences. Self-supervised pre-training provides LLMs with a broad understanding of textual and multimodal information, while supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) are essential for adapting these models to excel in targeted applications and adhere to user preferences and societal values. Data-centric SFT involves training models on carefully curated datasets across multiple modalities. In contrast, data-centric RLHF focuses on collecting human judgments to guide the model toward generating desirable and ethically aligned responses. Emphasizing data-centric approaches highlights the importance of high-quality, domain-specific, and multimodal datasets in the successful adaptation of MLLMs.